OpenStudy (anonymous):

find the interior angles of the triangle whose sides are 12cm,15cm,22cm. WITH MEDAL

OpenStudy (anonymous):

Do you know how to start?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

Have you heard of the Cosine Rule?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

It's sort of similar to the Pythagorean theorem but it works for all triangles. \[a^2 = b^2 +c^2 - 2(b)(c)\cos(\alpha)\]

OpenStudy (anonymous):



OpenStudy (anonymous):

So just plug those numbers into your calculator and solve.

OpenStudy (anonymous):

wait so the answer is

OpenStudy (anonymous):

its veryv hard :(

OpenStudy (anonymous):

It's ok, you can do it...

OpenStudy (anonymous):

teach me step by step

OpenStudy (anonymous):

Did you try plugging in those numbers into the formula I gave you?

OpenStudy (anonymous):

yes 12^2=15^2+22^2-2(15)(22) cos

OpenStudy (anonymous):

ok, now multiply those out

OpenStudy (anonymous):

144=225+484-660 cos

OpenStudy (anonymous):

144 = 225 + 484 - 660 cos(A) Now, simplify the expression so that you have "cos(A)" on one side and everything else on the other side.

OpenStudy (anonymous):

144=49cosA?

OpenStudy (anonymous):

No, more like this: \[\frac{ 144 -225 - 484 }{ -660 } = \cos(A) \]

OpenStudy (anonymous):

And then calculate the left side . . .

OpenStudy (anonymous):

+484

OpenStudy (anonymous):

What do you mean?

OpenStudy (anonymous):

144-225+484 over -660 =cosA

OpenStudy (anonymous):

No, because you are subtracting "225" AND "484" from both sides... Thus, the numerator should be "144 - 225 - 484"

OpenStudy (anonymous):

So, you are subtracting 225 from both sides and you are subtracting 484 from both sides.

OpenStudy (anonymous):

ok :)

OpenStudy (anonymous):

so the answer is 30.68

OpenStudy (anonymous):

You mean the angle?

OpenStudy (anonymous):

thats my final answer 30.68

OpenStudy (anonymous):

You mean the size of the angle in degrees?

OpenStudy (anonymous):

It's about that--I calculated that it was about 31.123 degrees... And remember, that's for angle A, the angle across from side "a", which we're assuming has length of 12 cm.

OpenStudy (anonymous):

Anyway, you continue the same process with another side like so: \[b^2 = a^2 + c^2 -2(a)(c)\cos(B)\] And then, when you have angle A & angle B, you can easily figure out the size of angle C because: C = 180 - A - B

OpenStudy (anonymous):

Make sense?

OpenStudy (anonymous):

@Splash_Dance how

OpenStudy (anonymous):

What?

OpenStudy (anonymous):

Splash_Dance

OpenStudy (anonymous):

Yes...

OpenStudy (anonymous):

interior angle?

OpenStudy (anonymous):

can i use herons law?

OpenStudy (anonymous):

No, that's for calculating the Area of a triangle--you're just looking for the angles.

OpenStudy (anonymous):